If P = 13%, x = 45, n = 250, then the computed z value is 2.351


1
Expert's answer
2021-06-10T05:17:10-0400
"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}"




"\hat{p}=\dfrac{x}{n}=\dfrac{45}{250}=0.18"


"p_0=0.13"

"z=\dfrac{0.18-0.13}{\sqrt{\dfrac{0.13(1-0.13)}{25}}}\approx2.351"

If P = 13%, x = 45, n = 250, then the computed z value is 2.351. True.